# Observable Optimization for Precision Theory: Machine Learning Energy Correlators

### v1

@@ 1 Introduction
As $k$ increases, $k$-point EECs capture in increasing detail the underlying physics at different energy scales. The simplest EEC observable, the two-point function, suffices to capture soft-collinear physics from QCD jets Kardos:2018kqj ; Gao:2019ojf ; Gao:2023ivm , and the two-pronged decays of the $W$ boson Ricci:2022htc . Analytic computations of multi-point EECs are currently limited to $3$ point in QCD Yang:2022tgm and $4$ point in $\mathcal{N}=4$ SYM Chicherin:2024ifn . In this work, we focus on using EECs to measure the top-quark mass, following Holguin:2022epo ; Holguin:2023bjf ; Holguin:2024tkz ; Xiao:2024rol . The top quark decays to three hard partons, so it is natural to search within the space of $3$-point EECs for sensitivity to the top-quark mass. In particular, we will look at highly energetic top quarks since the boost makes the three decay particles relatively collimated and the associated EEEC then becomes particularly well-suited for both measurement and precision theory. This was first considered in Ref. Holguin:2022epo , where the 1-dimension distribution of nearly equilateral-triangle three-point EECs was studied. More recent papers have looked at a couple of other shapes for three-point correlators Holguin:2023bjf ; Holguin:2024tkz ; Xiao:2024rol , finding increased performance compared to the equilateral case. This improvement motivates our systematic search for a marginalization which is optimally suited to top-quark mass measurement.
@@ 2.1 Simulation Details
Our study is performed on a dataset of hadronically-decaying top jets arising in $t\bar{t}$ events from $e^{+}e^{-}$ collisions at COM energy of $Q=2$ TeV that is generated with Pythia 8.309 Bierlich:2022pfr . The events were clustered using anti-$k_{t}$ jets of radius $R=1.2$ as implemented in Fastjet 3.4.1 Cacciari:2005hq . In order to reduce the large multiplicity in the resultant jets (the typical multiplicity of the anti-$k_{t}$ jets is around 150), the consituents were then reclustered via the Cambridge-Aachen algorithm Dokshitzer:1997in ; Wobisch:1998wt with radius $R^{\prime}=0.1$. This reduces the multiplicity to around 25. In addition, we placed a lower cutoff on the product of the energies of all triplets, selecting only events with $\frac{E_{i}E_{j}E_{k}}{Q^{3}}\geq 10^{-6}$. This cut significantly speeds processing and has little effect on the resulting EEECs. Then for each jet, we look at the roughly $\binom{25}{3}\sim 5000$ tuples and measure the angles and energies. This produces a distribution of points
@@ 2.3 DNN Density Approach
It is the second term which imposes the normalization constraint on the learnt function, and focuses the optimization over $L^{1}$ functions.
@@
Convergence of the loss function (assuming infinite statistics and perfect training) in Eq. (15) is guaranteed as it is based on the KL divergence between $\text{EEEC}_{\rm Data}$ and $\text{EEEC}_{\phi}$. Additionally, $\mathcal{L}_{\text{EEEC},\lambda}(\bm{\phi})$ is physically intuitive, since it not only prioritizes regions where samples are concentrated, but also re-weights them by their energy like the original observable. As with learning the ordinary probability distribution, the normalization term in Eq. (15) is needed to regularize the loss function and make the optimization problem well-posed.
@@ 2.3.2 Testing Marginals
The second marginalization we learn is of the EEEC as a function of a single side. In the left panel of Fig. 4, we show the marginalization in the large $\delta_{a}$ limit down to the largest side $\zeta_{\text{max}}=\zeta_{3}$, which includes all events (as in Chen:2020vvp ). In the right panel of Fig. 4, we plot a single example of the parameterization
@@ 2.4.1 Training Details
Our flow is a five dimensional, ten block deep flow implemented using nflowsnflows and modeled on the architecture from Krause:2021ilc ; Shih2022 . Each block consists of a rational quadratic spline (RQS) implemented using Masked Piecewise Rational Quadratic Autoregressive Transforms, a batch normalization layer, and a random permutation. For the RQS layers, we also use: ReLU activations, 40 bins, 200 hidden features, 2 context features, a tail bound of 14, min bin widths of $10^{-6}$, linear tails, and turn off residual blocks. [^2] We also tested flows with residual blocks as in Behrmann_invert_residual_net , but found they both slowed training and decreased performance. The base distribution is taken to be Gaussian in the four dimensions corresponding to energy and $\zeta$, and uniform between [-1,2] in the dimension corresponding to $m_{t}$ to avoid edge effects. The network was trained for 313 epochs (9 days on our gpu), where each epoch consists of reading 500,000 tuples randomly sampled from each mass (with tuples computed from the full 1M events), and each training batch consists of 500,000 tuples randomly mixed between masses. The flow is trained using the Adam optimizer with initial learning rate $10^{-4}$, with the learning rate reduced by a factor of 2 after 20 epochs without improvement, and early stopping after 50 epochs without improvement. We also tested energy weighting the flow loss, but it did not noticeably improve performance.
@@ 3.1 Architecture and Training
Figure 6: A schematic of the NRE architecture used to obtain the posterior $p(m_{t}|\vec{\zeta}_{\rm Shape})=p(m_{t})\exp(f_{\bm{\phi}}(\vec{\zeta}_{\rm Shape},m_{t}))$, including a summary network $s_{\bm{\phi}}$ for the shape $\vec{\zeta}_{\rm Shape}$ and a classifier network $c_{\phi}$.
@@
Table 2: Architecture details for the NRE network
@@
$\vec{\zeta}_{\text{shape}}=[\text{EEEC}(\zeta^{i=0}_{1},n_{1},n_{2},m_{t})_{\delta_{a}=0.01}(1+\epsilon_{i=0}),\text{EEEC}(\zeta^{i=1}_{1},n_{1},n_{2},m_{t})_{\delta_{a}=0.01}(1+\epsilon_{i=1}),\\ \cdots,\text{EEEC}(\zeta^{i=100}_{1},n_{1},n_{2},m_{t})_{\delta_{a}=0.01}(1+\epsilon_{i=100})]$
@@ 3.2 Comparing Shapes
While we expect our method for shape comparison to be robust, the actual number corresponding to this error should be taken with a grain of salt. The error here is primarily statistical, due to the statistical variation added to the density estimator during NRE training, and reflects the real statistical variation between shapes from data. That statistical noise is the source of error can be confirmed by seeing that the NRE error decreases if less noise is used during NRE training. In order to include other sources of error such as systematic errors which are represented by varying Monte Carlo parameters, an additional study explicitly including MC variations would need to be performed.
@@ 5 Acknowledgements
The authors thank Aurélien Dersy, Jesse Thaler, Rikab Gambhir, Ben Nachman, Vinicius Mikuni, Matthew Reece, Dennis Noll, Sascha Diefenbachar, and David Shih for useful discussions. AB and MDS are supported by DOE grant DE-SC0013607. KF is supported in part by: DOE grant DE-SC0013607, the Harvard GSAS Merit Fellowship, and the Miller Institute for Basic Research in Science, University of California Berkeley. KF also thanks the The Munich Institute for Astro-, Particle and BioPhysics and the Aspen Center for Physics (which is supported by NSF grant PHY-2210452, Simons Foundation grant (1161654, Troyer), and Alfred P. Sloan Foundation grant G-2024-22395) for hospitality while working on this project. The computations in this paper were performed on the Harvard Cannon Cluster, including resources provided by the Institute for Artificial Intelligence and Fundamental Interactions (IAIFI).
@@ Appendix A Classical Fits
For each, we computed the corresponding EEEC restricted such that for the smallest side $\zeta_{1}$, $3\zeta_{1}\in[0.02,0.1]$, and then fit a 15 degree polynomial to the histogram. We analytically found the local maximum for each and use the median $\zeta_{\mathrm{peak}}$ at each $m_{t}^{\text{MC}}$ to obtain the best fit $a$ and $b$ with Eq. (29). We then compute the standard deviation in $m_{t}$ values extracted using Eq. (30) and call this the statistical error for the linear fit. An example is shown in Fig. 10 for shapes $(n_{1},n_{2})_{\delta_{a}}=(1.02,1.99)_{0.01}$, and $(n_{1},n_{2})_{\delta_{a}}=(1.0,1.0)_{0.01}$. We find that the polynomial fit outlined above does yield errors that are of the same magnitude as that from the NRE, showing that statistical uncertainty dominates both the NRE and polynomial fit predictions for $m_{t}$. Because of this, we expect the NRE to be a relatively reliable proxy for the classical error, even though the correlation between the two is imperfect and is an interesting direction for future work. Additionally, we also check that the $m_{t}^{\text{MC}}$ value extracted using a linear fit with higher statistics (from full set of triplets from 1M events) typically lies within the NRE confidence interval, further improving our confidence in the NRE predictions. This is shown in Fig. 11a.

### v2

@@ 1 Introduction
As $k$ increases, $k$-point EECs capture in increasing detail the underlying physics at different energy scales. The simplest EEC observable, the two-point function, suffices to capture soft-collinear physics from QCD jets Kardos:2018kqj ; Gao:2019ojf ; Gao:2023ivm , and the two-pronged decays of the $W$ boson Ricci:2022htc . Analytic computations of multi-point EECs are currently limited to $3$ point in QCD Yang:2022tgm and $4$ point in $\mathcal{N}=4$ SYM Chicherin:2024ifn . In this work, we focus on using EECs to measure the top-quark mass, following Holguin:2022epo ; Holguin:2023bjf ; Holguin:2024tkz ; Xiao:2024rol . The top quark decays to three hard partons, so it is natural to search within the space of $3$-point EECs for sensitivity to the top-quark mass. In particular, we will look at highly energetic top quarks since the boost makes the three decay particles relatively collimated and the associated EEEC ($3$-point EEC) then becomes particularly well-suited for both measurement and precision theory. This was first considered in Ref. Holguin:2022epo , where the 1-dimension distribution of nearly equilateral-triangle three-point EECs was studied. More recent papers have looked at a couple of other shapes for three-point correlators Holguin:2023bjf ; Holguin:2024tkz ; Xiao:2024rol , finding increased performance compared to the equilateral case. This improvement motivates our systematic search for a marginalization which is optimally suited to top-quark mass measurement.
@@ 2.1 Simulation Details
Our study is performed on a dataset of hadronically-decaying top jets arising in $t\bar{t}$ events from $e^{+}e^{-}$ collisions at centre-of-mass energy of $Q=2$ TeV that is generated with Pythia 8.309 Bierlich:2022pfr . The events were clustered using anti-$k_{t}$ jets of radius $R=1.2$ as implemented in Fastjet 3.4.1 Cacciari:2005hq . In order to reduce the large multiplicity in the resultant jets (the typical multiplicity of the anti-$k_{t}$ jets is around 150), the consituents were then reclustered via the Cambridge-Aachen algorithm Dokshitzer:1997in ; Wobisch:1998wt with radius $R^{\prime}=0.1$. This reduces the multiplicity to around 25. In addition, we placed a lower cutoff on the product of the energies of each triplet, selecting only those with $\frac{E_{i}E_{j}E_{k}}{Q^{3}}\geq 10^{-6}$. This cut significantly speeds processing and has little effect on the resulting EEECs. Then for each jet, we look at the roughly $\binom{25}{3}\sim 5000$ tuples and measure the angles and energies. This produces a distribution of points
@@ 2.3 DNN Density Approach
It is the second term which imposes the normalization constraint on the learnt function, and focuses the optimization over $L^{1}$ functions. The discrete sum over $N$ and $M$ occurs over the batch used during training.
@@
Convergence of the loss function (assuming infinite statistics and perfect training) in Eq. (15) is guaranteed as it is based on the KL divergence between $\text{EEEC}_{\rm Data}$ and $\text{EEEC}_{\phi}$. Additionally, $\mathcal{L}_{\text{EEEC},\lambda}(\bm{\phi})$ is physically intuitive, since it not only prioritizes regions where samples are concentrated, but also re-weights them by their energy like the original observable. As with learning the ordinary probability distribution, the normalization term in Eq. (15) is needed to regularize the loss function and make the optimization problem well-posed. We remind the reader that the discrete sum over $N$ and $M$ occurs over the batch used for training the network
@@ 2.3.2 Testing Marginals
The second marginalization we evaluate is the EEEC as a function of a single side. In the left panel of Fig. 4, we show the marginalization in the large $\delta_{a}$ limit down to the largest side $\zeta_{\text{max}}=\zeta_{3}$, which includes all events (as in Chen:2020vvp ). In the right panel of Fig. 4, we plot a single example of the parameterization
@@ 2.4.1 Training Details
Our flow is a five dimensional, ten block deep flow implemented using nflowsnflows and modeled on the architecture from Krause:2021ilc ; Shih2022 . Each block consists of a rational quadratic spline (RQS) implemented using Masked Piecewise Rational Quadratic Autoregressive Transforms, a batch normalization layer, and a random permutation. For the RQS layers, we also use: ReLU activations, 40 bins, 200 hidden features, 2 context features, a tail bound of 14, min bin widths of $10^{-6}$, linear tails, and turn off residual blocks. [^2] We also tested flows with residual blocks as in Behrmann_invert_residual_net , but found they both slowed training and decreased performance. The base distribution is taken to be Gaussian in the four dimensions corresponding to energy and $\zeta$, and uniform between [1,2][-1,2] in the dimension corresponding to $m_{t}$ to avoid edge effects. The network was trained for 313 epochs (9 days on our GPU), where each epoch consists of reading 500,000 tuples randomly sampled from each mass (with tuples computed from the full 1M events), and each training batch consists of 500,000 tuples randomly mixed between masses. The flow is trained using the Adam optimizer with initial learning rate $10^{-4}$, with the learning rate reduced by a factor of 2 after 20 epochs without improvement, and early stopping after 50 epochs without improvement. We also tested energy weighting the flow loss, but it did not noticeably improve performance.
@@ 3.1 Architecture and Training
Figure 6: A schematic of the NRE architecture used to obtain the posterior $p(m_{t}|\vec{\zeta}_{\rm Shape})=p(m_{t})\exp(f_{\bm{\phi}}(\vec{\zeta}_{\rm Shape},m_{t}))$, including a summary network $s_{\bm{\phi}}$ for the shape $\vec{\zeta}_{\rm Shape}$ and a classifier network $c_{\phi}$. Gaussian noise $\epsilon$ at $5\%$ level was added to the density curves of the shapes to mimic histogram sampling noise, as described in Eqn.(23).
@@
Table 2: Architecture details for the NRE network. Both networks were trained simultaneously using a learning rate of $10^{-3}$ and a batch-size of 2000.
@@
$\vec{\zeta}_{\text{shape}}=[\text{EEEC}(\zeta^{i=0}_{1},n_{1},n_{2},m_{t})_{\delta_{a}=0.01}(1+\epsilon_{i=0}),\text{EEEC}(\zeta^{i=1}_{1},n_{1},n_{2},m_{t})_{\delta_{a}=0.01}(1+\epsilon_{i=1}),\\ \cdots,\text{EEEC}(\zeta^{i=100}_{1},n_{1},n_{2},m_{t})_{\delta_{a}=0.01}(1+\epsilon_{i=100})]$ (23)
@@ 3.2 Comparing Shapes
Thus, the grid search finds a shape (27) where the triangle is nearly isosceles in the small angle limit, with the ratio of largest to smallest side being approximately $\sqrt{n_{2}^{\star}}\sim\sqrt{2}$, (i.e) right isosceles triangles. While we expect our method for shape comparison to be robust, the actual number corresponding to this error should be taken with a grain of salt. The error here is primarily statistical, due to the statistical variation added to the density estimator during NRE training, and reflects the real statistical variation between shapes from data. That statistical noise is the source of error can be confirmed by seeing that the NRE error decreases when less noise is used during NRE training. Another potential source of error is the approximation of the DNN to the true EEEC distribution. We studied this in Fig. 8b by evaluating the NRE classifier, which was trained using the surrogate model, on pythia data, finding that the difference between the surrogate model and data is not an important source of error because the posterior for the true simulation is well within the envelope of posteriors for the DNN. In order to include other sources of error such as systematic errors which are represented by varying Monte Carlo parameters, an additional study explicitly including MC variations would need to be performed. One might also want to understand the contribution to the error from other sources, such as from the neural network itself. While important, there is currently no consensus on quantifying these sources of error and studying them is a significant open question in the field (see Bhimji:2024bcd ; Benato:2025rgo ; Khot:2025kqg ; Elsharkawy:2025six ; Ghosh:2021roe ; Nachman:2019dol ; Gambhir:2022gua ; Bollweg:2019skg ; Bellagente:2021yyh ; Rover:2024pvr ; Bieringer:2024nbc ; Bahl:2024gyt ; Benevedes:2025nzr for examples).
@@ 5 Acknowledgements
The authors thank Aurélien Dersy, Jesse Thaler, Rikab Gambhir, Ben Nachman, Vinicius Mikuni, Matthew Reece, Dennis Noll, Sascha Diefenbachar, and David Shih for useful discussions. AB and MDS are supported by DOE grant DE-SC0013607. KF is supported in part by: DOE grant DE-SC0013607, the Harvard GSAS Merit Fellowship, and the Miller Institute for Basic Research in Science, University of California Berkeley. This work is also supported by the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/). KF also thanks the The Munich Institute for Astro-, Particle and BioPhysics and the Aspen Center for Physics (which is supported by NSF grant PHY-2210452, Simons Foundation grant (1161654, Troyer), and Alfred P. Sloan Foundation grant G-2024-22395) for hospitality while working on this project. The computations in this paper were performed on the Harvard Cannon Cluster, including resources provided by the Institute for Artificial Intelligence and Fundamental Interactions (IAIFI).
@@ Appendix A Classical Fits
For each, we computed the corresponding EEEC restricted such that for the smallest side $\zeta_{1}$, $3\zeta_{1}\in[0.02,0.1]$, and then fit a 15 degree polynomial to the histogram. We analytically found the local maximum for each and use the median $\zeta_{\mathrm{peak}}$ at each $m_{t}^{\text{MC}}$ to obtain the best fit $a$ and $b$ with Eq. (30). We then compute the standard deviation in $m_{t}$ values extracted using Eq. (31) and call this the statistical error for the linear fit. An example is shown in Fig. 10 for shapes $(n_{1},n_{2})_{\delta_{a}}=(1.02,1.99)_{0.01}$, and $(n_{1},n_{2})_{\delta_{a}}=(1.0,1.0)_{0.01}$. We find that the polynomial fit outlined above does yield errors that are of the same magnitude as that from the NRE, showing that statistical uncertainty dominates both the NRE and polynomial fit predictions for $m_{t}$. Because of this, we expect the NRE to be a relatively reliable proxy for the classical error, even though the correlation between the two is imperfect and is an interesting direction for future work. Additionally, we also check that the $m_{t}^{\text{MC}}$ value extracted using a linear fit with higher statistics (from full set of triplets from 1M events) typically lies within the NRE confidence interval, further improving our confidence in the NRE predictions. This is shown in Fig. 11.